# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v6

@@ Abstract
Using the semiformal method in combinatorics we generalize Pólya’s theorem. This theorem determines limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$ starting at the origin visit the given vertex. We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$. In part II we consider analogs of Pólya’s theorem for edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
Besides the abstract, we want to help the reader understand our article by the following preamble. In it we have more space to explain what we do.
We mention Pólya’s theorem, which is a well known result in the discipline of random walks. However, our article is not really about generalizing Pólya’s theorem. We use extensions of Pólya’s theorem to edge weighted countable complete graphs merely as a test case for the novel generalization of the well known ([3, 8, 10, 12, 22, 23, 28]) symbolic method in enumerative (analytic) combinatorics. This generalization is due to the first author, and we call it the semiformal method. It will be developed in detail in [14]. The present article is an explanation of the functioning of the semiformal method in the form of an extensive example. More applications of it are given in [15]. The idea of the semiformal method is to extend finite sums defining coefficients of generating functions in the symbolic method to infinite (countable) series. Finite sums in a ring (in the symbolic method, one also needs to multiply) are straightforward, but it is less clear what is a sum of an infinite series of real or complex numbers. We use three kinds of sums of infinite series: absolute convergence (which is used in probability where summands are nonnegative), the more general and weaker conditional convergence, and sums equal to $\pm\infty$. The last sums are in fact important — the heart of the combinatorial proof of the case $d\leq 2$ in Pólya’s theorem is the fact that a certain series has sum $\infty$. For details on the uses of these three kinds of sums we refer to the article below. But why confine ourselves to infinite series of real or complex numbers, when non-Archimedean infinite series are also interesting? This is the topic of [17]. To summarize, the main contribution of this article is the demonstration of the working of the semiformal method in the fourteen cases that arise when Pólya’s theorem is generalized to $K_{\mathbb{N}}$ with complex edge weights.
@@ 1 Introduction
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$ starting at $\overline{0}$ visits in a step $i>0$ the vertex $\overline{v}$. In 1921, G. Pólya proved in [24] a theorem, which we generalize, that for every vertex $\overline{v}$,
@@
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\,.\end{array}\right$.
@@
Pólya’s theorem is almost always mentioned in the special case with $\overline{v}=\overline{0}$, but he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a quote where he states his result). A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 27, 31, 32].
We generalize it as follows. Let
$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\text{ with }|e|=2\}$}
be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$. Here $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) denotes the cardinality of a finite set $A$. Let
$h\colon\mathbb{N}_{2}\to\mathbb{C}$
be an edge weight, where $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$. Let $n\in\mathbb{N}_{0}$. We denote by $W(n)$ the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$. Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable. Let $v\in\mathbb{N}$. We denote by $W(v,n)$ ($\subset W(n)$) the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$ in a step $i>0$. We extend the weight multiplicatively to
$h\colon\bigcup_{n=0}^{\infty}W(n)\to\mathbb{C}\,.$}
Let $W\subset\bigcup_{n=0}^{\infty}W(n)$. If the series
$\sum_{w\in W}h(w)$}
absolutely converges, we denote its sum by $h(W)$. (Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges. Then so does the series $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
$A_{h,v}(x)=\sum_{n\geq 0}h(W(v,\,n))x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
Our generalizations of Pólya’s theorem are formulas for three quantities:
$\lim_{x\nearrow 1}A_{h,v}(x),\ A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,$,}
where $\lim_{x\nearrow 1}$ means that $x$ goes to $1$ through values $x\in[0,1)$. In some cases, these quantities are twisted by powers $\xi^{n}$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$). The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.34.10 in Section 4. Pólya’s theorem is subsumed in five of them. Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$. Theorems 4.7 (4.8) and 4.10 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
The fourteen generalizations come from three sources. First, weights in Pólya’s theorem are nonnegative real numbers, and the convergence of infinite series is absolute. For our complex weights, we consider two kinds of convergence: absolute and conditional. We also consider real series with sums $\pm\infty$. Second, Pólya’s theorem is on probabilities, which means that the weights of edges incident to a fixed vertex sum up to $\xi=1$. We generalize it to $\xi$-convexity for any $\xi\in\mathbb{C}$; the case of $\xi\neq 0$ is normalized to $|\xi|=1$ and the case $\xi=0$ is treated separately (it corresponds to Kirchhoff’s first law). Thus we consider two cases, general weights and $\xi$-convex weights. Third, Pólya’s theorem is, in reality, two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs. In our generalization, we correspondingly distinguish between cases $v=1$ and $v\neq 1$. Now $3\cdot 2\cdot 2=12$ and two additional generalizations arise by varying the assumptions.
We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we denote by
$f[C]=\{f(x)\colon\;x\in A\cap C\}\ \ (\subset B)$

### v8

# Extending the symbolic method in enumerative combinatorics. I

@@ Abstract
We use our extension of the symbolic method in enumerative combinatorics (we extend finite sums determining coefficients in generating functions to infinite series) to generalize Pólya’s theorem. This theorem finds the limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$, starting at the origin, visit the given vertex. We replace the grid graph with weighted countable complete graphs $K_{\mathbb{N}}$ such that the edge weights have values in $\mathbb{C}$ and satisfy a convergence condition. Thus general Pólya’s theorems are results in enumerative combinatorics.
@@ 1 Introduction
We recall in general terms the fundamental symbolic method in enumerative combinatorics, briefly the symbolic method, and explain how we extend it. Later in the article, we exemplify our extension on Pólya’s theorem. Our selection of references on the symbolic method is [3, 8, 11, 13, 29].
1.1 The symbolic method and its extension
Using this method, we enumerate—by exact counting formulas and by approximate asymptotic formulas—an infinite (countable) set $A$ of combinatorial objects. $A$ is equipped with a size function $s\colon A\to\mathbb{N}_{0}$ ($\{0,1,\dots\}$). It is assumed that for every $n\in\mathbb{N}_{0}$ the set of size $n$ objects in $A$,
$A_{n}=\{a\in A\colon\;s(a)=n\}\,$,
is finite; see [3, p. 1], [8, p. ix], [11, Definition I.1 on p. 16] and [13, Remark 2.2.3 on p. 32, implicitly]. We associate with $A$ and $s$ the generating function
$A(x)=\sum_{a\in A}x^{s(a)}=\sum_{n\geq 0}|A_{n}|x^{n}\ \ (\in\mathbb{C}[[x]])\,$,}
where $|X|$ denotes the cardinality of a finite set $X$. It is common, and crucial in the approach here, to have a weight function $h\colon A\to R$, where $R$ is a ring. The weighted version of $A(x)$ is then
$A_{h}(x)=\sum_{n\geq 0}(\sum_{a\in A_{n}}h(a))x^{n}\ \ (\in R[[x]])\,.$}
Note that every coefficient $\sum_{a\in A_{n}}h(a)$ in $A_{h}(x)$ is a finite sum. The symbolic method works in two phases. The first phase translates combinatorial relations between objects in $A$ into algebraic, differential, or functional equations for $A_{h}(x)$ and the auxiliary generating functions. Multivariate generating functions can also be used. The second phase deduces from these equations asymptotic formulas for the coefficients of generating functions. The second analytic phase accounts for the usefulness and popularity of the method (see [11]), but the first formal phase is the primary one (the memoir [3] is an impressive automatization of this phase).
We give two examples to illustrate the symbolic method: one short now and another, detailed in the next subsection. Let $A$ be the set of all proper bracketings $b$, let $s(b)$ be the number of pairs of brackets in $b$ and let $h(b)=1$. For example,
$A_{3}=\{()()(),\,()(()),\,(())(),\,(()()),\,((()))\}\,.$
The numbers $(|A_{n}|)_{n\geq 0}=(1,1,2,5,14,42,\dots)$ are the Catalan numbers [28]. The two phases of the symbolic method yield the following.
  1. 1.
    We consider the first pair of brackets in $b$ and get the algebraic equation $A(x)=1+xA(x)^{2}$ (we omit the index h because the weight is trivial). It leads to the following exact counting formulas.
    1. (a)
      $|A_{0}|=1$ and $|A_{n+1}|=\sum_{j=0}^{n}|A_{j}|\cdot|A_{n-j}|$ for $n\in\mathbb{N}_{0}$.
    2. (b)
      $|A_{n}|=\frac{1}{n+1}\binom{2n}{n}$ for $n\in\mathbb{N}_{0}$.
    3. (c)
      $|A_{0}|=1$ and $|A_{n+1}|=\frac{4n+2}{n+2}\cdot|A_{n}|$ for $n\in\mathbb{N}_{0}$.
    For $m\in\mathbb{N}_{0}$, let $v(m)=\log(m+2)$; up to multiplicative constant, $v(m)$ is the number of binary (or $q$-ary, for any $q\geq 2$) digits in $m$. Each of (a)–(c) is an efficient closed formula for the counting function $C:\;n\mapsto|A_{n}|$ because it provides an algorithm that computes $C$ in
    $O(\max(v(n),\,v(|A_{n}|))^{c})=O((n+1)^{c}),\ n\in\mathbb{N}_{0},\,$
    ($c>0$ is a constant) bit operations. See the next phase for the asymptotics of $|A_{n}|$.
  2. 2.
    By [11, p. 38], the equation $xA(x)^{2}-A(x)+1=0$ (or (b)) produces the approximate asymptotic formula
    $|A_{n}|=(1+o(1))\pi^{-1/2}n^{-3/2}4^{n}\ \ (n\to\infty)\,.$}
    See [11, Chapter VII.7] for the theory of asymptotics of coefficients of algebraic formal power series.
At the end of phase 1 we remind the definition of an efficient closed formula for a counting function, which was given by the first author in [15, p. 10, Wilfian formulas]. This example of application of the symbolic method is standard, maybe even archetypal. Our next example in the next subsection is less standard
Our idea of extension of the symbolic method is simple and natural. We drop the assumption that the sets $A_{n}$ are finite, and replace it with the assumption that the ring $R$ is endowed with a complete norm, for example, $R=\mathbb{R}$ or $R=\mathbb{C}$, and that all possibly infinite series
$\sum_{a\in A_{n}}h(a)$}
determining coefficients in $A_{h}(x)$ absolutely converge. This extension relies on our semiformal approach to formal power series. In this approach, we allow operations with formal power series in which coefficients arise from limit transitions. In this article, $R=\mathbb{C}$ and we work out an extensive example in which we apply the extended symbolic method to generalize Pólya’s theorem, as explained below. In [17] we develop the semiformal approach in general. In [19] we obtain analogs of the results in this article for the ring $R=\mathbb{C}[[x_{1},\dots,x_{k}]]$ which is endowed with a complete non-Archimedean norm. In [16] we use the semiformal approach in number theory and present two proofs of the transcendence of the number $\mathrm{e}$ based on formal power series.
1.2 Pólya’s theorem for $\overline{v}=\overline{0}$ via the symbolic method
@@
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$, starting at the origin $\overline{0}=\langle 0,0,\dots,0\rangle$ ($\in\mathbb{Z}^{d}$), later visits the vertex $\overline{v}$. In 1921, G. Pólya proved in [24] a theorem, which we generalize in our article, that for every dimension $d\in\mathbb{N}$ and every vertex $\overline{v}$,
@@
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ \in(0,\,1)&\dots&\text{if $d\geq 3$}\,.\end{array}\right$. (1)
@@
Pólya’s theorem is almost always mentioned in the special case when $\overline{v}=\overline{0}$, but Pólya proved it for every vertex $\overline{v}$ (see [18, Section 1] for a quote in which he states his result). A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 10, 12, 14, 20, 21, 22, 23, 25, 26, 27, 32, 33].